OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the following. sin 25Ï€/21 cos 6Ï€/7 - cos 25Ï€/21 sin 6Ï€/7

OpenStudy (anonymous):

could someone please explain?

OpenStudy (welshfella):

i think you can use the followingidentity here sin(A - B) =sinAcosB - cosAsinB

OpenStudy (anonymous):

So it would be sin 25 - 6?

OpenStudy (welshfella):

No sin(25Pi/21 - 6pi/7) = sin(25pi/21 - 18pi/21) can you ccontinue?

OpenStudy (anonymous):

Okay I got decimal though

OpenStudy (anonymous):

0.018

OpenStudy (welshfella):

you need to find the angle in terms of pi it = sin(7pi/21) = sin(pi/3) now using a standard triangle you can find the exact value of sin(pi/3)

OpenStudy (welshfella):